In a triangle, the measure of the second angle is 20 more than twice the first and the measure of the third angle is 55 more than one-third of the first. Find the measure of the angles of the triangle.
Answer by rfer(10417)

x=31.5 degrees
2x+20=83 degrees
1/3x+55=65.5 degrees
x+2x+20+1/3x+55=180
3 1/3x=105
x=105/(10/3)
x=105*(3/10)
x=315/10
x=31.5